OpenStudy (geometry_hater):

PLEASE HELP Find the surface area of a cone with a diameter of 13 and a slant height of 5. (use 3.14 for pi)

OpenStudy (anonymous):

r=6.5,l=5 surface area= pi r l + pi r^2 = pi r(r+l) now, substitute value of each

OpenStudy (geometry_hater):

whered 6.5 come from

OpenStudy (anonymous):

radius= half of diameter = 13/2 =6.5